I need help with this problem.
"As an orange grows over a five-week period, there is a linear relationship between the volume of the orange and time. At the beginning of the period, the volume is about 34cm^3. four weeks later, the volume is about 50cm^3. What is the volume at the end of the five-week period?"
Please help!! Thank you!!!

Answer by stanbon(75887)

I need help with this problem.
"As an orange grows over a five-week period, there is a linear relationship between the volume of the orange and time. At the beginning of the period, the volume is about 34cm^3. four weeks later, the volume is about 50cm^3. ---------------------------------------
Your have two points: (0,34) and (4,50)
----
Linear Form: V = mw + b where v is volume and w is weeks.
---
Find "m" = slope:
m = (50-34)/4 = 4
-----
intercept = V(0) = 34
-----
Equation:
V(w) = 4w+34
-----
What is the volume at the end of the five-week period?"
V(5) = 4*5+34
V(5) = 54 cm^3
===================
Cheers,
Stan H.
===================